# Towards Human-AI Complementarity with Predictions Sets

### v1

@@ 6 Experiments with Real Data
Figure 3: Complementary cumulative distribution (cCDF) of the per-image test accuracy achieved by a simulated human expert using the prediction sets constructed with conformal prediction (Naive and Aps), our greedy algorithm (Greedy), and brute force search (Brute Force Search) on the ImageNet16H dataset.
Results. Table 2 and Figure 3 show the average test accuracy and complementary cumulative distribution (cCDF) of the test per-image accuracy achieved by a simulated human expert using the prediction sets constructed with conformal prediction (Naive and APS), our greedy algorithm (Greedy) and brute force search (Brute Force Search) for different values of noise $\omega$. The results show that, similarly as in our experiments with synthetic data, the greedy algorithm achieves the same performance as brute force search. Moreover, they also show that, using greedy algorithm to construct prediction sets, the expert achieves the highest average accuracy in all group of images except the group with $\omega=80$, where both the greedy algorithm and one of the conformal predictors offer comparable performance.
7 Discussion and Limitations
Hardness analysis. In our hardness analysis, our reduction utilizes an instance of our problem in which, for every prediction set, the predictions made by experts follow a mixture of MNLs. As an immediate consequence, this implies that, in general, the problem of finding the prediction set ${\mathcal{S}}$ that maximizes $\hat{g}({\mathcal{S}}\,|\,x)$ is NP-hard to approximate. However, in our experiments, the greedy algorithm is almost always able to find such a set ${\mathcal{S}}$. As a result, we hypothesize that there may be certain conditions on the parameters of the mixture of MNLs under which the problem can be efficiently approximated to a factor less than the size of the label set.
Evaluation. The results of our experiments suggest that the prediction sets constructed using our greedy algorithm may help human experts make more accurate predictions than the prediction sets constructed using conformal prediction. However, one may argue that the difference in performance is partly due to the fact that the non-conformity scores used in conformal prediction do not incorporate information about the distribution of experts’ predictions. Motivated by this observation, it would be important to investigate how to incorporate such information in the definition of non-conformity scores. Moreover, in our experiments, the true distribution of experts’ predictions matches the mixture of MNLs used by the greedy algorithm and brute force search. However, in practice, there may a mismatch between the true distribution of experts’ predictions and the mixture of MNLs, and this may decrease performance.
@@ 8 Conclusions
We have looked at the problem of finding the optimal prediction sets under which human experts achieve the highest accuracy in a given multiclass classification task. We have shown that this problem is NP-hard to solve and to approximate to any factor less than the size of the label set. However, we have empirically shown that, for a large parameterized class of expert models, a simple greedy algorithm consistently finds (near-)optimal prediction sets offering greater performance than those constructed using conformal prediction.
Acknowledgements. Gomez-Rodriguez acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 945719).
@@ A.2 Proof of Theorem 1
and $\widehat{N}_{{\mathcal{S}}}(y)$ denotes the number of vertices that are not adjacent to $y$ in the subgraph induced [^13] A graph $\mathcal{G}_{\mathcal{S}}=({\mathcal{S}},\mathcal{E}_{\mathcal{S}})$ is a vertex-induced subgraph in graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, if ${\mathcal{S}}\subseteq\mathcal{V}$ and for every $u,v\in{\mathcal{S}}$, $(u,v)\in\mathcal{E}_{\mathcal{S}}$ if and only if $(u,v)\in\mathcal{E}$. $A_{{\mathcal{S}}}(u)$ denote the set of vertices adjacent to $u$ in $\mathcal{G}_{\mathcal{S}}$ and $\widehat{A}_{{\mathcal{S}}}(u)={\mathcal{S}}\setminus A_{{\mathcal{S}}}(u)$ denote the set of vertices that are not adjacent to $u$. Note that, for any $u\in{\mathcal{S}}$, we assume $u\notin A_{{\mathcal{S}}}(u)$ that implies $u\in\widehat{A}_{{\mathcal{S}}}(u)$. Similarly, $N_{{\mathcal{S}}}(u)=|A_{{\mathcal{S}}}(u)|$ and $\widehat{N}_{{\mathcal{S}}}(u)=|\widehat{A}_{{\mathcal{S}}}(u)|$ denote the corresponding subset sizes. by ${\mathcal{S}}$. The transformation described above is dominated by the size of $\bm{C}$, which is $O(|\mathcal{V}|^{2})$, so the reduction is polynomial time.
@@ Appendix B Running time analysis of the greedy algorithm
In this section, we present the complexity analysis for the greedy algorithm in Algorithm 1. Adding each element requires computing the marginal gain $\Delta=\hat{g}({\mathcal{S}}_{k}\cup\{y\}\,|\,x)-\hat{g}({\mathcal{S}}_{k}\,|\,x)$ at most $k$ times, where $k$ is the size of ${\mathcal{S}}_{k}$. This computation in Line 8 can be efficiently performed in $O(k)$ time. Specifically, it can be rewritten as:
$\hat{g}({\mathcal{S}}_{k}\cup\{y\}\,|\,x)-\hat{g}({\mathcal{S}}_{k}\,|\,x)=\sum_{\hat{y}\in{\mathcal{S}}_{k}\cup\{y\}}f_{\hat{y}}(x)\frac{C_{\hat{y}\hat{y}}}{\underbrace{\sum_{y^{\prime}\in{\mathcal{S}}_{k}\cup\{y\}}C_{y^{\prime}\hat{y}}}_{(a)}}-\sum_{\hat{y}\in{\mathcal{S}}_{k}}f_{\hat{y}}(x)\frac{C_{\hat{y}\hat{y}}}{\sum_{y^{\prime}\in{\mathcal{S}}_{k}}C_{y^{\prime}\hat{y}}}.$
The term $(a)$ (i.e., $\sum_{y^{\prime}\in{\mathcal{S}}_{k}\cup\{y\}}C_{y^{\prime}\hat{y}}$) in the denominator can be computed as $\sum_{y^{\prime}\in{\mathcal{S}}_{k}}C_{y^{\prime}\hat{y}}+C_{y\hat{y}}$ by storing the value of $\sum_{y^{\prime}\in{\mathcal{S}}_{k}}C_{y^{\prime}\hat{y}}$ at the end of each iteration after Line 11. The loops in Line 5 and Line 7 each iterate over $k$, resulting in $O(k^{3})$ iterations for each value of $k\in\{1,\dots,L\}$. So, the overall running time of our algorithm is $O(L^{4})$.
@@ Appendix C Implementation details for the experiments with synthetic and real data
Experiments with synthetic data. We employ the make_classification utility function of scikit-learn to generate various classification tasks. It is a convenient method to generate $L$-class classification tasks by varying several parameters such as the task difficulty, the number of labels and the number of informative features. It generates $n$ clusters of points positioned on the vertices of a $d$-dimensional hypercube by adding interdependencies and noise to the features. In Section 5, we set the number of features to 20, the number of redundant features to 0 and the number of informative features to $d=4$ (for a $L=10$ label classification task). We assign a balanced proportion of samples for each class. We control the task difficulty by choosing the class_sep parameter, which represents the length of the sides of the hypercubes, thus indicating how far apart are the various classes. A smaller class_sep implies a more difficult classification task. We vary the class_sep parameter to ensure the classifier and the humans span different ranges of accuracies. Please refer to the original documentation for more information. [^15] https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_classification.html#sklearn.datasets.make_classification
@@ Appendix D Empirical average test accuracy on additional classification tasks
Table 3: Empirical average test accuracy achieved by four different (simulated) human experts, each with a different noise value $\gamma$, on their own (None) and using prediction sets constructed with conformal prediction (Naive, Aps) and with the greedy algorithm (Greedy) on four synthetic classification tasks. In each table, we vary the number of labels $L\in\{25,50\}$. Each cell shows the average and standard deviation over $10$ runs. We denote the best results for each classification task in bold.
@@
Table 5: (b) $L=50$
Appendix E Empirical conditional probability that a prediction set includes $\{y,\bar{y}\}$ given $Y=y$ for additional classification tasks

### v2

@@ 7 Discussion and Limitations
Figure 3: Complementary cumulative distribution (cCDF) of the per-image test accuracy achieved by a simulated human expert using the prediction sets constructed with conformal prediction (Naive, Aps, Raps and Saps) and our greedy algorithm (Greedy) on the ImageNet16H dataset.
@@ 8 Conclusions
We have looked at the problem of finding the optimal prediction sets under which human experts achieve the highest accuracy in a given multiclass classification task. We have shown that this problem is $\mathsf{NP}$-hard to solve and to approximate to any factor less than the size of the label set. However, we have empirically shown that, for a large parameterized class of expert models, a simple greedy algorithm consistently outperforms conformal prediction.
Acknowledgements. Gomez-Rodriguez acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 945719). De Toni acknowledges support from the TANGO project (grant #101120763-TANGO). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the EU or HaDEA. Neither the EU nor the granting authority can be held responsible for them.
@@ A.2 Proof of Theorem 1
and $\widehat{N}_{{\mathcal{S}}}(y)$ denotes the number of vertices that are not adjacent to $y$ in the subgraph induced [^15] A graph $\mathcal{G}_{\mathcal{S}}=({\mathcal{S}},\mathcal{E}_{\mathcal{S}})$ is a vertex-induced subgraph in graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, if ${\mathcal{S}}\subseteq\mathcal{V}$ and for every $u,v\in{\mathcal{S}}$, $(u,v)\in\mathcal{E}_{\mathcal{S}}$ if and only if $(u,v)\in\mathcal{E}$. by ${\mathcal{S}}$. The transformation described above is dominated by the size of $\bm{C}$, which is $O(|\mathcal{V}|^{2})$, so the reduction is polynomial time.
@@ Appendix B Running time analysis of the greedy algorithm
In this section, we present the complexity analysis for the greedy algorithm in Algorithm 1. Adding each element requires computing the marginal gain $\Delta=\hat{g}({\mathcal{S}}\cup\{y\}\,|\,x)-\hat{g}({\mathcal{S}}\,|\,x)$ at most $k$ times. This computation in Line 8 can be efficiently performed in $O(k)$ time, where $k$ is the size of ${\mathcal{S}}_{k}$. Specifically, it can be rewritten as:
$\hat{g}({\mathcal{S}}_{k}\cup\{y\}\,|\,x)-\hat{g}({\mathcal{S}}_{k}\,|\,x)=\sum_{\hat{y}\in{\mathcal{S}}_{k}\cup\{y\}}f_{\hat{y}}(x)\frac{C_{\hat{y}\hat{y}}}{\sum_{y^{\prime}\in{\mathcal{S}}_{k}\cup\{y\}}C_{y^{\prime}\hat{y}}}-\sum_{\hat{y}\in{\mathcal{S}}_{k}}f_{\hat{y}}(x)\frac{C_{\hat{y}\hat{y}}}{\sum_{y^{\prime}\in{\mathcal{S}}_{k}}C_{y^{\prime}\hat{y}}}.$
The term $\sum_{y^{\prime}\in{\mathcal{S}}_{k}\cup\{y\}}C_{y^{\prime}\hat{y}}$ in the denominator can be computed as $\sum_{y^{\prime}\in{\mathcal{S}}_{k}}C_{y^{\prime}\hat{y}}+C_{y\hat{y}}$ by storing the value of $\sum_{y^{\prime}\in{\mathcal{S}}_{k}}C_{y^{\prime}\hat{y}}$ at the end of each iteration after Line 11. The loops in Line 5 and Line 7 each iterate over $k$, resulting in $O(k^{3})$ iterations for each value of $k\in\{1,\dots,L\}$. So, the overall running time of our algorithm is $O(L^{4})$.
@@ Appendix C Implementation details for the experiments with synthetic and real data
Experiments with synthetic data. We employ the make_classification utility function of scikit-learn to generate the various prediction task. It is a convenient method to generate $L$-class classification tasks by varying several parameters such as the task difficulty, the number of labels and the number of informative features. It generates $n$ clusters of points positioned on the vertices of a $d$-dimensional hypercube by adding interdependencies and noise to the features. In Section 5, we set the number of features to 20, the number of redundant features to 0 and the number of informative features to $d=4$ (for a $L=10$ label classification task). We assign a balanced proportion of samples for each class. We control the task difficulty by choosing the class_sep parameter, which represents the length of the sides of the hypercubes, thus indicating how far apart are the various classes. A smaller class_sep implies a more difficult classification task. We vary the class_sep parameter to ensure the classifier and the humans span different ranges of accuracies. Please refer to the original documentation for more information. [^17] https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_classification.html#sklearn.datasets.make_classification
@@ Appendix D Empirical average test accuracy on additional classification tasks
In this section, we keep the same experimental setting as described in Section 5 but we vary the number of labels $L\in\{25,50\}$ of the classification task. Table D summarizes the results.
Table 3: Empirical average test accuracy for $L\in\{25,50\}$.
@@
Raps $0.395\,\scriptstyle\pm 0.011$ $0.601\,\scriptstyle\pm 0.008$ $0.874\,\scriptstyle\pm 0.009$ $0.939\,\scriptstyle\pm 0.008$
Saps $0.396\,\scriptstyle\pm 0.010$ $0.601\,\scriptstyle\pm 0.008$ $0.876\,\scriptstyle\pm 0.013$ $0.943\,\scriptstyle\pm 0.007$
@@
Raps $0.380\,\scriptstyle\pm 0.010$ $0.585\,\scriptstyle\pm 0.009$ $0.863\,\scriptstyle\pm 0.011$ $0.933\,\scriptstyle\pm 0.008$
Saps $0.380\,\scriptstyle\pm 0.009$ $0.585\,\scriptstyle\pm 0.011$ $0.863\,\scriptstyle\pm 0.014$ $0.937\,\scriptstyle\pm 0.007$
@@
Raps $0.363\,\scriptstyle\pm 0.009$ $0.560\,\scriptstyle\pm 0.011$ $0.841\,\scriptstyle\pm 0.011$ $0.920\,\scriptstyle\pm 0.007$
Saps $0.362\,\scriptstyle\pm 0.009$ $0.558\,\scriptstyle\pm 0.015$ $0.840\,\scriptstyle\pm 0.014$ $0.920\,\scriptstyle\pm 0.008$
@@
Raps $0.346\,\scriptstyle\pm 0.011$ $0.529\,\scriptstyle\pm 0.016$ $0.823\,\scriptstyle\pm 0.013$ $0.907\,\scriptstyle\pm 0.009$
Saps $0.344\,\scriptstyle\pm 0.010$ $0.528\,\scriptstyle\pm 0.017$ $0.822\,\scriptstyle\pm 0.014$ $0.907\,\scriptstyle\pm 0.009$
@@
Raps $0.412\,\scriptstyle\pm 0.017$ $0.625\,\scriptstyle\pm 0.016$ $0.791\,\scriptstyle\pm 0.013$ $0.932\,\scriptstyle\pm 0.009$
Saps $0.425\,\scriptstyle\pm 0.020$ $0.635\,\scriptstyle\pm 0.034$ $0.802\,\scriptstyle\pm 0.014$ $0.934\,\scriptstyle\pm 0.009$
@@
Raps $0.398\,\scriptstyle\pm 0.018$ $0.612\,\scriptstyle\pm 0.016$ $0.781\,\scriptstyle\pm 0.014$ $0.927\,\scriptstyle\pm 0.010$
Saps $0.407\,\scriptstyle\pm 0.021$ $0.620\,\scriptstyle\pm 0.033$ $0.793\,\scriptstyle\pm 0.014$ $0.931\,\scriptstyle\pm 0.009$
@@
Raps $0.380\,\scriptstyle\pm 0.010$ $0.587\,\scriptstyle\pm 0.016$ $0.754\,\scriptstyle\pm 0.015$ $0.916\,\scriptstyle\pm 0.012$
Saps $0.387\,\scriptstyle\pm 0.008$ $0.590\,\scriptstyle\pm 0.029$ $0.764\,\scriptstyle\pm 0.013$ $0.920\,\scriptstyle\pm 0.009$
@@
Raps $0.367\,\scriptstyle\pm 0.014$ $0.560\,\scriptstyle\pm 0.015$ $0.722\,\scriptstyle\pm 0.017$ $0.906\,\scriptstyle\pm 0.012$
Saps $0.367\,\scriptstyle\pm 0.017$ $0.556\,\scriptstyle\pm 0.028$ $0.725\,\scriptstyle\pm 0.020$ $0.907\,\scriptstyle\pm 0.012$
@@
Table 5: $L=50$
Appendix E Empirical average test coverage of prediction sets across classification tasks
In this section, we report the empirical average test coverage across synthetic and real classification tasks achieved by the prediction sets constructed using conformal prediction and our greedy algorithms. Table 6 and Table 7 summarizes the results, which show that the prediction sets constructed using the best conformal predictors and our greedy algorithm achieve comparable empirical coverage. Moreover, in those few settings in which the prediction sets constructed using conformal prediction achieve higher coverage (e.g., $P(Y^{\prime}=Y)=0.3$ and $\gamma=0.3$), the average test accuracy achieved by the simulated human experts using the prediction sets constructed using conformal prediction is lower (cf. Table 1), which underlines how coverage alone may be a bad proxy for estimating the average accuracy achieved by human experts using prediction sets.
Table 6: Empirical average test coverage achieved by the prediction sets constructed using conformal prediction (Naive, Aps, Raps, Saps) and using the greedy algorithm (Greedy) on four synthetic classification tasks with four different (simulated) human experts, each with a different noise value $\gamma$. For each classification task, the classifier $f$ used by conformal prediction and the greedy algorithm achieves a different average accuracy $P(Y^{\prime}=Y)$. For each (simulated) human expert, the best value of $\alpha$ for each conformal predictor is different and thus the reported coverage is different. The number of labels is $L=10$ and, the size of the calibration set is $m=1000$. Each cell shows the average and standard deviation over $10$ runs. We denote the best results for each task in bold.
$\gamma$ Method $\mathbb{P}[Y^{\prime}=Y]=0.3$ $\mathbb{P}[Y^{\prime}=Y]=0.5$ $\mathbb{P}[Y^{\prime}=Y]=0.7$ $\mathbb{P}[Y^{\prime}=Y]=0.9$
0.3 Naive $\bm{0.637\,\scriptstyle\pm 0.066}$ $0.802\,\scriptstyle\pm 0.058$ $0.908\,\scriptstyle\pm 0.020$ $0.973\,\scriptstyle\pm 0.007$
Aps $0.603\,\scriptstyle\pm 0.083$ $\bm{0.804\,\scriptstyle\pm 0.045}$ $0.900\,\scriptstyle\pm 0.026$ $0.967\,\scriptstyle\pm 0.006$
Raps $0.592\,\scriptstyle\pm 0.087$ $0.792\,\scriptstyle\pm 0.032$ $0.911\,\scriptstyle\pm 0.015$ $0.965\,\scriptstyle\pm 0.012$
Saps $0.580\,\scriptstyle\pm 0.067$ $0.794\,\scriptstyle\pm 0.041$ $0.890\,\scriptstyle\pm 0.013$ $0.965\,\scriptstyle\pm 0.010$
Greedy $0.502\,\scriptstyle\pm 0.024$ $0.764\,\scriptstyle\pm 0.019$ $\bm{0.920\,\scriptstyle\pm 0.012}$ $\bm{0.976\,\scriptstyle\pm 0.004}$
0.5 Naive $\bm{0.583\,\scriptstyle\pm 0.104}$ $\bm{0.770\,\scriptstyle\pm 0.044}$ $0.897\,\scriptstyle\pm 0.021$ $0.968\,\scriptstyle\pm 0.009$
Aps $0.557\,\scriptstyle\pm 0.100$ $0.741\,\scriptstyle\pm 0.036$ $0.879\,\scriptstyle\pm 0.016$ $0.961\,\scriptstyle\pm 0.007$
Raps $0.534\,\scriptstyle\pm 0.096$ $0.761\,\scriptstyle\pm 0.040$ $0.844\,\scriptstyle\pm 0.043$ $0.950\,\scriptstyle\pm 0.012$
Saps $0.557\,\scriptstyle\pm 0.081$ $0.768\,\scriptstyle\pm 0.042$ $0.876\,\scriptstyle\pm 0.014$ $0.961\,\scriptstyle\pm 0.009$
Greedy $0.489\,\scriptstyle\pm 0.027$ $0.732\,\scriptstyle\pm 0.015$ $\bm{0.902\,\scriptstyle\pm 0.013}$ $\bm{0.970\,\scriptstyle\pm 0.003}$
0.7 Naive $\bm{0.535\,\scriptstyle\pm 0.104}$ $0.676\,\scriptstyle\pm 0.047$ $0.823\,\scriptstyle\pm 0.044$ $0.938\,\scriptstyle\pm 0.013$
Aps $0.519\,\scriptstyle\pm 0.115$ $0.661\,\scriptstyle\pm 0.036$ $0.853\,\scriptstyle\pm 0.015$ $0.941\,\scriptstyle\pm 0.013$
Raps $0.506\,\scriptstyle\pm 0.071$ $0.644\,\scriptstyle\pm 0.056$ $0.813\,\scriptstyle\pm 0.019$ $0.938\,\scriptstyle\pm 0.010$
Saps $0.492\,\scriptstyle\pm 0.079$ $\bm{0.721\,\scriptstyle\pm 0.051}$ $0.858\,\scriptstyle\pm 0.031$ $0.942\,\scriptstyle\pm 0.011$
Greedy $0.473\,\scriptstyle\pm 0.017$ $0.696\,\scriptstyle\pm 0.014$ $\bm{0.861\,\scriptstyle\pm 0.014}$ $\bm{0.958\,\scriptstyle\pm 0.005}$
1.0 Naive $\bm{0.499\,\scriptstyle\pm 0.075}$ $0.608\,\scriptstyle\pm 0.034$ $0.750\,\scriptstyle\pm 0.025$ $0.905\,\scriptstyle\pm 0.014$
Aps $0.453\,\scriptstyle\pm 0.062$ $0.631\,\scriptstyle\pm 0.038$ $0.806\,\scriptstyle\pm 0.026$ $0.912\,\scriptstyle\pm 0.013$
Raps $0.466\,\scriptstyle\pm 0.071$ $0.611\,\scriptstyle\pm 0.024$ $0.787\,\scriptstyle\pm 0.026$ $0.920\,\scriptstyle\pm 0.013$
Saps $0.484\,\scriptstyle\pm 0.070$ $0.609\,\scriptstyle\pm 0.063$ $0.764\,\scriptstyle\pm 0.022$ $0.907\,\scriptstyle\pm 0.006$
Greedy $0.457\,\scriptstyle\pm 0.024$ $\bm{0.664\,\scriptstyle\pm 0.015}$ $\bm{0.839\,\scriptstyle\pm 0.014}$ $\bm{0.951\,\scriptstyle\pm 0.006}$
Table 7: Empirical coverage achieved by the prediction sets constructed using conformal prediction (Naive, Aps, Raps, Saps) and using the greedy algorithm (Greedy) on the ImageNet-16H dataset. Each cell shows the average and standard deviation over $10$ runs. We denote the best results for each noise level in bold.
Method $\omega=80$ $\omega=95$ $\omega=110$ $\omega=125$
Naive $\bm{0.977\,\scriptstyle\pm 0.005}$ $\bm{0.972\,\scriptstyle\pm 0.009}$ $\bm{0.962\,\scriptstyle\pm 0.010}$ $0.923\,\scriptstyle\pm 0.018$
Aps $0.970\,\scriptstyle\pm 0.009$ $0.962\,\scriptstyle\pm 0.007$ $0.943\,\scriptstyle\pm 0.015$ $0.887\,\scriptstyle\pm 0.009$
Raps $0.967\,\scriptstyle\pm 0.009$ $0.956\,\scriptstyle\pm 0.007$ $0.944\,\scriptstyle\pm 0.013$ $0.876\,\scriptstyle\pm 0.015$
Saps $0.967\,\scriptstyle\pm 0.009$ $0.966\,\scriptstyle\pm 0.009$ $0.946\,\scriptstyle\pm 0.008$ $0.921\,\scriptstyle\pm 0.011$
Greedy $0.975\,\scriptstyle\pm 0.008$ $\bm{0.972\,\scriptstyle\pm 0.010}$ $\bm{0.962\,\scriptstyle\pm 0.008}$ $\bm{0.931\,\scriptstyle\pm 0.007}$
Appendix F Empirical conditional probability that a prediction set includes $\{y,\bar{y}\}$ given a ground-truth label $Y=y$ for additional classification tasks